# CITRAS: Covariate-Informed Transformer for Time Series Forecasting

### v1

@@ B.1 Full Results of Multivariate forecasting
For a comprehensive comparison, we evaluate our model on well-recognized long-term multivariate forecasting datasets. We adopt the rolling forecasting approach [23], where a model trained with predetermined input and forecasting lengths is evaluated with a longer forecasting length by iteratively integrating its output into the subsequent input. Following previous works [23, 24], we set the input length to 672, the forecasting length to {96, 192, 336, 720}, and the patch length to 96. The full results are listed in Table 5.
@@
Table 5: Full results of the multivariate forecasting, where the input length is set to 672, and the forecasting length is set to {96, 192, 336, 720}. Results are cited from Timer-XL [24] if available; otherwise reproduced. #$1^{\text{st}}$ denotes the first count, and #$1^{\text{st}},2^{\text{nd}}$ denotes the sum of the first and second counts.
@@
TFT
Cross.
TimesNet
Auto.

### v2

@@ F.2 Full Results of Multivariate Forecasting
The full results of multivariate forecasting in Section 4.2 are listed in Table 10. We adopt the rolling forecasting approach [29], where a model trained with predetermined input and forecasting lengths is evaluated with a longer forecasting length by iteratively integrating its output into the subsequent input. Following previous works [29, 30], we set the input length to 672, the training forecasting length to 96, the testing forecasting length to {96, 192, 336, 720}, and the patch length to 96.
@@
Table 10: Full results of the multivariate forecasting, where the input length is set to 672, and the forecasting length is set to {96, 192, 336, 720}. The evaluation protocol is the same as in Table 2. Results are cited from Timer-XL [30] if available; otherwise reproduced.
@@
Leddam
CARD
ModernTCN
FITS
DLinear